Ray BD bisects angle ABC. The measure of angle ABD is represented by 4x-4 and the measure of angle ABC is represented by 7x+4. Find the measures of angle DBC
Answer #1

image from custom entry tool

Since BD bisetcts angle ABC, the measure of angle ABD = the measure of angle DBC = 4x-4 = 1/2 of measure of angle ABC =

->

-> x = 12

so the measure of angle DBC = 4*12-4 = 44

answered by: Jessie! :)